OpenStudy (anonymous):

I REALLY NEED HELP Mike poured 15 cups of water into a rectangular container measuring 10 inches by 6 inches by 9 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. Part B: How many cubic inches of water was poured into the container? Part C: What was the height of the water in the container? Show your work. Measurement is given as length by width by height.

Directrix (directrix):

The volume of the rectangular container is given by the formula: V = L*W*H V = 10" * 6" * 9 cm V = 10 * 6 * 22.86 (22.86 = 9 * 2.54 to change the 9 cm to inches) V = 1371.6 cubic inches @LegendaryNikki Please check this and let me know what you think, okay?

Directrix (directrix):

@LegendaryNikki I'm waiting for you.

OpenStudy (anonymous):

I dont know what to do @Directrix

Directrix (directrix):

You could start by checking the volume work above. @LegendaryNikki

OpenStudy (anonymous):

I dont know how to check it thats how bad i am at algebra.. @Directrix

Directrix (directrix):

B) Mike poured 15 cups of water into the container. 1 cup = 14.44 cubic inches @LegendaryNikki --Here's something you can do --Multiply 14.44 times 15 and post what you get.

Directrix (directrix):

@LegendaryNikki Post this product so we can finish this --> -Multiply 14.44 times 15 and post what you get.

OpenStudy (anonymous):

216.6

Directrix (directrix):

Okay, so 216.6 cubic inches is the answer to this: Part B: How many cubic inches of water was poured into the container?

OpenStudy (anonymous):

okay

Directrix (directrix):

And, this may be the answer to Part A if you check it: V = L*W*H V = 10" * 6" * 9 cm V = 10 * 6 * 22.86 (22.86 = 9 * 2.54 to change the 9 cm to inches) V = 1371.6 cubic inches

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

The pan's capacity is V = 1371.6 cubic inches. Mike poured 216.6 cubic inches into the pan.

Directrix (directrix):

216.6 = 10 * 6 * h 216.6 = 60 h h = 216.6 / 60 h = ? @LegendaryNikki --> Please divide 216.6 by 60 and post what you get.

OpenStudy (anonymous):

3.61

Directrix (directrix):

Okay, then the height of the water Mike poured into the pan is 3.61 inches. So, that's it, I think.

OpenStudy (anonymous):

Thank you @Directrix

Directrix (directrix):

You are welcome.

Directrix (directrix):

@LegendaryNikki

OpenStudy (lena772):

@Directrix You're supposed to divide by 2.54 to find the inches.